# ThermalGaussian: Thermal 3D Gaussian Splatting

### v1

@@ Abstract
Thermography is especially valuable for the military and other users of surveillance cameras. Some recent methods based on Neural Radiance Fields (NeRF) are proposed to reconstruct the thermal scenes in 3D from a set of thermal and RGB images. However, unlike NeRF, 3D Gaussian splatting (3DGS) prevails due to its rapid training and real-time rendering. In this work, we propose ThermalGaussian, the first thermal 3DGS approach capable of rendering high-quality images in RGB and thermal modalities. We first calibrate the RGB camera and the thermal camera to ensure that both modalities are accurately aligned. Subsequently, we use the registered images to learn the multimodal 3D Gaussians. To prevent the overfitting of any single modality, we introduce several multimodal regularization constraints. We also develop smoothing constraints tailored to the physical characteristics of the thermal modality. Besides, we contribute a real-world dataset named RGBT-Scenes, captured by a hand-hold thermal-infrared camera, facilitating future research on thermal scene reconstruction. We conduct comprehensive experiments to show that ThermalGaussian achieves photorealistic rendering of thermal images and improves the rendering quality of RGB images. With the proposed multimodal regularization constraints, we also reduced the model’s storage cost by 90%. The code and dataset will be released.
@@
Figure 1: Compared to previous NeRF-based methods (Hassan et al., 2024) and approaches that directly use thermal images for training 3DGS (Kerbl et al., 2023), our approach not only achieves the highest thermal image rendering quality but also significantly reduces the model’s storage requirements through multimodal regularization (MR).
@@ 1 Introduction
Previous 3D thermal scene reconstruction (Rangel et al., 2014; Zhao et al., 2017; Müller, 2019; Li et al., 2023) typically involves a two-stage process. In the first stage, RGB images and traditional multi-view geometry methods (Newcombe et al., 2011) are used to achieve a 3D geometric reconstruction of the scene. In the second stage, thermal images are mapped onto the reconstructed 3D scene. However, these methods not only fail to fully exploit thermal information but are also constrained by the limitations of traditional 3D reconstruction techniques, which impede their ability to render high-quality images from a new view. This significantly hinders the practical application of thermal reconstruction.
@@
(2)We propose a series of strategies for multimodal Gaussian reconstruction, including multimodal initialization, three different thermal Gaussians training strategies, constraints specific to thermal modalities, and multimodal regularization.
@@ 2.1 Thermal imaging and 3D reconstruction
All objects with temperatures above absolute zero emit energy in the form of electromagnetic waves, a phenomenon known as thermal radiation. Through Planck’s law, thermal imaging converts the wavelength and intensity of electromagnetic waves radiated from an object’s surface into thermal information, which is then used to create images. Originally, thermal imaging was developed for military purposes to enable visualization under extreme lighting conditions, such as nighttime or smoke. As costs have decreased in recent years, it has also been widely applied in fields of healthcare, industry, agriculture, building inspection, and so on.
3D reconstruction has significant applications in visualizing, surveying, and analyzing large, inaccessible, and important objects, in ways that 2D images cannot provide. The advent of the groundbreaking KinectFusion (Newcombe et al., 2011) marked the beginning of an era of high-precision, dense 3D reconstruction. Subsequent developments (Nießner et al., 2013; Kähler et al., 2015; Dai et al., 2017; Gong et al., 2021; Zhang et al., 2021) have optimized 3D reconstruction in terms of accuracy, efficiency, and unconstrained camera movement. Some works (Rangel et al., 2014; Zhao et al., 2017; Müller, 2019; Li et al., 2023) have been made to integrate thermal imaging with the aforementioned 3D reconstruction methods, leading to the development of thermal 3D reconstruction. However, these traditional multi-view geometry-based methods do not perform as well in rendering new views as the more recent deep learning-based approaches.
NeRF (Mildenhall et al., 2021) has emerged as a significant milestone in the field of 3D reconstruction due to its impressive ability to render highly realistic images from a new view. Recently, ThermoNeRF (Hassan et al., 2024) and Thermal-NeRF (Ye et al., 2024) have been proposed to reconstruct thermal scenes by combining a set of thermal images with NeRF. Although these approaches successfully generate images from new perspectives, they are limited by the slow rendering speed and implicit scene representation of NeRF, which hampers their practical application.
@@
3DGS (Kerbl et al., 2023) represent a revolutionary technology in the fields of 3D reconstruction. Distinct from methods like NeRF, 3DGS employs millions of explicit Gaussians, fundamentally altering its approach. This technology merges the advantages of neural network-based optimization with structured data representation, enabling photorealistic rendering from new views, significantly enhancing real-time rendering capabilities, and introducing the ability to manipulate and edit 3D scenes. These features make 3DGS highly compatible with a broad range of downstream applications, establishing it as the baseline for next-generation 3D reconstruction technologies (Chen & Wang, 2024). Although 3DGS is constrained and trained using only RGB modality images, it ultimately generates millions of Gaussians, resembling a point cloud. This characteristic makes 3DGS particularly suitable for multimodal fusion with other devices that directly capture scene point clouds, such as depth cameras and LiDAR. Studies (Matsuki et al., 2024; Yan et al., 2024; Keetha et al., 2024) have effectively integrated depth cameras to implement 3DGS-based simultaneous localization and mapping. Studies (Li et al., 2024; Chung et al., 2024) have effectively combined depth images (Yan et al., 2020) estimated from a pre-trained Monocular Depth estimation model (Bhat et al., 2023) with the RGB modality, resulting in improved rendering quality and more accurate geometric structures. 3D scenes contain not only RGB and geometric modalities but also other important modalities relevant to various applications, such as thermal, material, and pressure. We integrate 3DGS with thermal modality.
@@ 3.2 Multimodal Initialization
Previously, methods for calibration RGB and thermal images (Zhang et al., 2023) often involve designing specialized, non-standard metallic calibration boards with uniformly sized circular holes, as shown in Fig. 4a. The calibration relies on the temperature difference between the board and the background to compute thermal features, enabling calibration. However, the high complexity and stringent requirements for producing these calibration boards make them difficult to obtain and lack a universal standard. We find that a standard chessboard pattern, as shown in Fig 4b, commonly used for RGB camera calibration, can effectively be used for calibrating both thermal and color cameras, with a mean reprojection error of less than 0.5 pixels. Initially, we heat the calibration board using devices like an infrared heater; black regions, absorbing heat faster due to their material properties, exhibit relatively higher temperatures. We capture simultaneous color and thermal images before a thermal equilibrium with surroundings is achieved. Subsequently, conventional camera calibration (Zhang, 1999) is performed.
@@ 3.3 Thermal Gaussian
Multiple Single-Modal Gaussians (MSMG): The training of MFTG initially utilizes RGB modal information followed by thermal modal information. Although both modalities are employed, they are not used simultaneously. This sequential approach during the thermal Gaussian training in the second phase may not fully leverage the information from the color modality. Therefore, in MSMG (as shown in Fig. 3), we constrain the training with information from both color and thermal modalities simultaneously. We train two single-modal Gaussians initialized by point clouds from multimodal initialization. The thermal Gaussian renders thermal images, while the RGB Gaussian renders RGB images. Subsequently, these rendered images of both modalities are compared separately with the ground truth of their respective inputs using loss functions:
@@
One Multi-Modal Gaussian(OMMG): OMMG extends MSMG by not only employing dual-modal loss constraints but also integrating multiple modalities onto a single Gaussian. This integration ensures that information from diverse modalities is unified within a single geometric structure. Specifically, we construct a multimodal Gaussian comprising positional coordinates $x$, scaling matrix $\mathbf{S}$, rotation matrix $\mathbf{R}$, opacity $\alpha$, spherical harmonics $c$ for RGB representation, and spherical harmonics $t$ for thermal representation. RGB rendering is achieved using Equation 3, while thermal rendering follows the equation below:
@@
3.4 Thermal Loss and Multimodal Regularization
@@
When training multiple modalities of the same object simultaneously, it is often undesirable for one modality to dominate at the expense of others. Therefore, a regularization strategy is needed to dynamically adjust the weight of each modality’s loss during training. In the training of MSMG, we observe that the weight coefficients of a modality align linearly with the number of generated Gaussian. A higher weight for a modality results in more Gaussian generated by that modality, while fewer Gaussian are generated by another modality. Hence, we design multi-modal regularization coefficients $\gamma$ based on the number of Gaussian generated by each modality during training.
@@ 5.1 Implementation Details
Our method is an improvement upon the 3DGS framework, with all experimental settings consistent with the reference 3DGS. The specific hyperparameter $\lambda_{smooth}$ is set to 0.6. Each comparative experiment was trained for 30K iterations. All experiments are conducted on a single NVIDIA 3090 GPU. The resolution of the rendered RGB images and thermal images is 640×480.
@@ 5.2 Thermal View Synthesis
Similar to 3DGS, we employ image quality assessment metrics including Peak Signal-to-Noise Ratio (PSNR) (Hore & Ziou, 2010), Structural Similarity Metric (SSIM) (Wang et al., 2004), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) to evaluate the quality of reconstructed thermal and RGB images from new views.
@@ 5.3 RGB View Synthesis
Figure 6: We present qualitative RGB image comparisons between our method and 3DGS from test views.
Our method not only achieves high-quality thermal image rendering but also significantly enhances RGB image rendering quality. As shown quantitatively in Table 4, our multimodal constraints improve RGB rendering quality in nearly all scenarios, with an average PSNR improvement of 1.1 dB compared to the original 3DGS (Kerbl et al., 2023). This improvement is particularly evident in scenarios where the RGB modality struggles to identify the environment, while the thermal modality can recognize it clearly. As shown in the top part of Fig. 6, where distinguishing between foreground and background is challenging in the RGB modality but straightforward in the thermal modality due to temperature differences, constraints from the thermal modality aid in the accurate learning of the RGB modality. Additionally, as depicted in the bottom part of Fig. 6, the assistance from thermal images enables accurate color rendering in low-light scenes for the RGB modality. Our results demonstrate that, under multimodal constraints, when one modality fails, our approach leverages accurate information from the other modality to enhance the model’s understanding of the scene, thus facilitating the correct learning of the failing modality. This enables our method to advance 3D reconstruction in low-light scenes and enhances the robustness of 3D reconstruction techniques to some extent.
@@
+ $\mathcal{L}_{\text{smooth }}$ /  - 24.65 0.867 0.198 - - - 603s 292 61MB
OursMFTG ✓  /  - 24.70 0.871 0.191 - - - 491s+575s 316 51MB
@@ 5.4 Ablation Study
We separate different contributions and algorithm choices to test their effectiveness. As shown in Table 2 and Table 4, after incorporating multimodal initialization, allows 3DGS to achieve thermal reconstruction across various environments. Our designed multimodal thermal Gaussian models, MSMG and OMMG, not only render both thermal and RGB images simultaneously but also improve rendering quality for both modalities in all scenes, with an average increase of over 1.2 dB. We also observed that multimodal constraints mitigate the generation of excessive redundant Gaussians in certain modalities. Later, we introduced a regularization term to dynamically adjust the coefficients of both modalities. As shown in Table 4, directly training RGB modality Gaussians with 3DGS results in an average storage requirement of 159 MB. On the other hand, directly training thermal Gaussians with MI requires an average of 65 MB. The RGB Gaussians for OursMSMG+MR average only 9 MB in storage, with thermal Gaussians averaging the same. Our method requires only 8% $(\frac{9+9}{159+65}=0.08)$ of the storage space compared to directly using 3DGS. Due to the reduction in the number of Gaussians, the rendering speed has also significantly increased. Additionally, the rendering quality for both modalities has also improved. In Fig. 7a, we compare our multimodal regularization $\gamma$ with manually adjusting the thermal constraint coefficients in a truck scene. The comparison shows that our multimodal regularization approach reduces storage space for both RGB and thermal modalities while maintaining high image quality. In Fig.7b, we visually present the Gaussian distributions of the original 3DGS method and our method with multimodal regularization.
@@ 6 Conclusion and Future Work
We are the first to implement thermal reconstruction based on 3DGS. We not only achieve simultaneous rendering of thermal and RGB images but also significantly improve the rendering quality of both color and thermal images. Additionally, we greatly reduce the model’s storage requirements. In the future, we plan to incorporate super-resolution for more detailed thermal reconstruction and to explore the reconstruction of dynamic thermal processes.

### v2

@@ Abstract
Thermography is especially valuable for the military and other users of surveillance cameras. Some recent methods based on Neural Radiance Fields (NeRF) are proposed to reconstruct the thermal scenes in 3D from a set of thermal and RGB images. However, unlike NeRF, 3D Gaussian splatting (3DGS) prevails due to its rapid training and real-time rendering. In this work, we propose ThermalGaussian, the first thermal 3DGS approach capable of rendering high-quality images in RGB and thermal modalities. We first calibrate the RGB camera and the thermal camera to ensure that both modalities are accurately aligned. Subsequently, we use the registered images to learn the multimodal 3D Gaussians. To prevent the overfitting of any single modality, we introduce several multimodal regularization constraints. We also develop smoothing constraints tailored to the physical characteristics of the thermal modality. Besides, we contribute a real-world dataset named RGBT-Scenes, captured by a hand-hold thermal-infrared camera, facilitating future research on thermal scene reconstruction. We conduct comprehensive experiments to show that ThermalGaussian achieves photorealistic rendering of thermal images and improves the rendering quality of RGB images. With the proposed multimodal regularization constraints, we also reduced the model’s storage cost by 90%. Our project page is at https://thermalgaussian.github.io/.
@@
Figure 1: Compared to NeRF-based methods (Hassan et al., 2024) and methods that directly use thermal images for training 3DGS, our approach not only improves the thermal image rendering quality but also significantly reduces the model’s storage through multimodal regularization (MR).
@@ 1 Introduction
Previous 3D thermal scene reconstruction (Rangel et al., 2014; Zhao et al., 2017; Müller, 2019; Li et al., 2023) involves a two-stage process. In the first stage, RGB images and traditional multi-view geometry methods (Newcombe et al., 2011) are used to achieve a 3D geometric reconstruction. In the second stage, thermal images are mapped onto the reconstructed 3D scene. However, these methods not only fail to fully exploit thermal information but are also constrained by the limitations of traditional 3D reconstruction techniques, which impede their ability to render high-quality images from a new view. This hinders the practical application of thermal reconstruction.
@@
(2)We propose a series of strategies for multimodal Gaussian reconstruction, including multimodal initialization, three different thermal Gaussians, constraints specific to thermal modalities, and multimodal regularization.
@@ 2.1 Thermal imaging and 3D reconstruction
All objects with temperatures above absolute zero emit energy in the form of electromagnetic waves, a phenomenon known as thermal radiation. Through Planck’s law, thermal imaging converts the wavelength and intensity of electromagnetic waves radiated from an object’s surface into thermal information, which is then used to create images. Common thermal imaging devices operate in the mid-to-long-wave infrared range. To intuitively display the temperature distribution in thermal imaging, pseudo-color rendering is often applied using a cool-to-warm color scheme. Originally, thermal imaging was developed for military purposes to enable visualization under extreme lighting conditions, such as nighttime or smoke. As costs have decreased in recent years, it has also been widely applied in fields of healthcare, industry, agriculture, building inspection, and so on.
By capturing 2D images from various angles and applying 3D reconstruction, a 3D model can be created. Unlike fixed 2D viewpoints, the 3D model enables intuitive analysis from any perspective and scale, allowing detailed exploration. The advent of KinectFusion (Newcombe et al., 2011) marked the beginning of an era of high-precision, dense 3D reconstruction. Subsequent developments (Nießner et al., 2013; Kähler et al., 2015; Dai et al., 2017; Gong et al., 2021; Zhang et al., 2021) have optimized 3D reconstruction in terms of accuracy, efficiency, and unconstrained camera movement. Some works (Rangel et al., 2014; Zhao et al., 2017; Müller, 2019; Li et al., 2023) have been made to integrate thermal imaging with the aforementioned methods, leading to the development of thermal 3D reconstruction. However, these traditional multi-view geometry-based methods do not perform as well in rendering new views as the more recent deep learning-based approaches.
With the rapid development of artificial intelligence(Yan et al., 2020a; Wang et al., 2024a), NeRF (Mildenhall et al., 2021) has emerged as a significant milestone in the field of 3D reconstruction due to its impressive ability to render highly realistic images from a new view. Recently, ThermoNeRF (Hassan et al., 2024) and Thermal-NeRF (Ye et al., 2024) have been proposed to reconstruct thermal scenes by combining a set of thermal images with NeRF. Although these approaches successfully generate images from new perspectives, they are limited by the slow rendering speed and implicit scene representation of NeRF, which hampers their practical application.
@@
3DGS (Kerbl et al., 2023) represent a revolutionary technology in the fields of 3D reconstruction. Distinct from methods like NeRF, 3DGS employs millions of explicit Gaussians, fundamentally altering its approach. This technology merges the advantages of neural network-based optimization with structured data representation, enabling photorealistic rendering from new views, significantly enhancing real-time rendering capabilities, and introducing the ability to manipulate and edit 3D scenes. These features make 3DGS highly compatible with a broad range of downstream applications, establishing it as the baseline for next-generation 3D reconstruction technologies (Chen & Wang, 2024). Although 3DGS is constrained and trained using only RGB modality images, it ultimately generates millions of Gaussians, resembling a point cloud. This characteristic makes 3DGS particularly suitable for multimodal fusion with other devices that directly capture scene point clouds, such as depth cameras and LiDAR. Studies (Matsuki et al., 2024; Yan et al., 2024; Keetha et al., 2024) have effectively integrated depth cameras to implement 3DGS-based simultaneous localization and mapping. Studies (Li et al., 2024; Chung et al., 2024) have effectively combined depth images (Yan et al., 2020b) estimated from a pre-trained Monocular Depth estimation model (Bhat et al., 2023) with the RGB modality, resulting in improved rendering quality and more accurate geometric structures. 3D scenes contain not only RGB and geometric modalities but also other important modalities relevant to various applications, such as thermal, material, and pressure.
@@ 3.2 Multimodal Initialization
Previously, methods for calibration RGB and thermal images (Zhang et al., 2023) often involve designing specialized, non-standard metallic calibration boards with uniformly sized circular holes, as shown in Fig. 4a. The calibration relies on the temperature difference between the board and the background to compute thermal features, enabling calibration. However, the high complexity and stringent requirements for producing these calibration boards make them difficult to obtain and lack a universal standard. We find that a standard chessboard pattern, as shown in Fig 4b, commonly used for RGB camera calibration, can effectively be used for calibrating both thermal and color cameras, with a mean reprojection error of less than 0.5 pixels. Initially, we heat the calibration board using devices like an infrared heater; black regions, absorbing heat faster due to their material properties, exhibit relatively higher temperatures. We capture simultaneous color and thermal images before thermal equilibrium, which occurs when two systems reach a balanced state with equal temperatures, halting heat flow. Subsequently, conventional camera calibration (Zhang, 1999) is performed.
@@ 3.3 Thermal Gaussian
Multiple Single-Modal Gaussians (MSMG): The training of MFTG initially utilizes RGB modal information followed by thermal modal information. Although both modalities are employed, they are not used simultaneously. Since only thermal images were utilized for supervision in stage two, the information from the color modality was not fully leveraged. Therefore, in MSMG (as shown in Fig. 3), we constrain the training with information from both color and thermal modalities simultaneously. We train two single-modal Gaussians initialized by point clouds from multimodal initialization. The thermal Gaussian renders thermal images, while the RGB Gaussian renders RGB images. Subsequently, these rendered images of both modalities are compared separately with the ground truth of their respective inputs using loss functions:
@@
One Multi-Modal Gaussian(OMMG): OMMG extends MSMG by not only employing dual-modal loss constraints in Eq. (7) but also integrating multiple modalities onto a single Gaussian. This integration ensures that information from diverse modalities is unified within a single geometric structure. Specifically, we construct a multimodal Gaussian comprising positional coordinates $x$, scaling matrix $\mathbf{S}$, rotation matrix $\mathbf{R}$, opacity $\alpha$, spherical harmonics $c$ for RGB representation, and spherical harmonics $t$ for thermal representation. RGB rendering is achieved using Formula 3, while thermal rendering follows the equation below:
@@
3.4 Thermal Loss &\&& Multimodal Regularization
@@
When training multiple modalities of the same object simultaneously, it is often undesirable for one modality to dominate at the expense of others. Therefore, a regularization strategy is needed to dynamically adjust the weight of each modality’s loss during training. In the training of MSMG, we observe that the weight coefficients of a modality align linearly with the Gaussian it ultimately generates. A higher weight for a modality results in more Gaussian generated by that modality, while fewer Gaussian are generated by another modality. Hence, we design multi-modal regularization coefficients $\gamma$ based on the number of Gaussian generated by each modality during training.
@@ 5.1 Implementation Details
Our method is an improvement upon the 3DGS framework, with all experimental settings (e.g., $\lambda$) remaining consistent with the reference 3DGS. The specific hyperparameter $\lambda_{smooth}$ is set to 0.6. Each comparative experiment was trained for 30K iterations. All experiments are conducted on a single NVIDIA 3090 GPU. The resolution of the rendered RGB images and thermal images is 640×480.
@@ 5.2 Thermal View Synthesis
Similar to 3DGS, we employ image quality assessment metrics including Peak Signal-to-Noise Ratio (PSNR) (Hore & Ziou, 2010; Wang et al., 2024b), Structural Similarity Metric (SSIM) (Wang et al., 2004), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) to evaluate the quality of reconstructed thermal and RGB images from new views.
@@ 5.3 RGB View Synthesis
Figure 6: We present qualitative RGB image comparisons between our method and 3DGS.
Our method not only achieves high-quality thermal image rendering but also significantly enhances RGB image rendering quality. As shown quantitatively in Table 4, our multimodal constraints improve RGB rendering quality in nearly all scenarios, with an average PSNR improvement of 1.1 dB compared to the original 3DGS. This improvement is particularly evident in scenarios where the RGB modality struggles to identify the environment, while the thermal modality can recognize it clearly. As shown in the top of Fig.6, where distinguishing between foreground and background is challenging in the RGB modality but straightforward in the thermal modality due to temperature differences, constraints from the thermal modality aid in the accurate learning of the RGB modality. Additionally, as depicted in the bottom of Fig.6, the assistance from thermal images enables accurate color rendering in low-light scenes for the RGB modality. Our results demonstrate that, under multimodal constraints, when one modality fails, our approach leverages accurate information from the other modality to enhance the model’s understanding of the scene, thus facilitating the correct learning of the failing modality. This enables our method to advance 3D reconstruction in low-light scenes and enhances the robustness of 3D reconstruction techniques to some extent.
@@
+MI+$\mathcal{L}_{\text{smooth }}$ /  - 24.65 0.867 0.198 - - - 603s 292 61MB
OursMFTG ✓  /  - 24.70 0.871 0.191 - - - 491s 316 51MB
@@ 5.4 Ablation Study
We separate different contributions and algorithm choices to test their effectiveness. As shown in Table 2 and Table 4, after incorporating multimodal initialization, allows 3DGS to achieve thermal reconstruction across various environments. Our multimodal thermal Gaussian models, MSMG and OMMG, not only render both thermal and RGB images simultaneously but also improve rendering quality for both modalities in all scenes, with an average increase of over 1.2 dB. We also observed that multimodal constraints mitigate the generation of excessive redundant Gaussians. Later, we introduced a regularization term to dynamically adjust the coefficients of both modalities. As shown in Table 4, directly training RGB modality Gaussians with 3DGS results in an average storage requirement of 159 MB. On the other hand, directly training thermal Gaussians with MI requires an average of 65 MB. The RGB Gaussians for MSMG+MR average only 9 MB in storage, with thermal Gaussians averaging the same. Our method requires only 8% $(\frac{9+9}{159+65}=0.08)$ of the storage space compared to directly using 3DGS. Due to the reduction in the number of Gaussians, the rendering speed has also significantly increased. Additionally, the rendering quality for both modalities has also improved. MFTG, MSMG+MR, and OMMG excel in different aspects: training speed, storage efficiency, and rendering quality. In Fig.7 a, we compare our multimodal regularization $\gamma$ with manually adjusting the thermal constraint coefficients in a truck scene. The comparison shows that our multimodal regularization approach reduces storage space for both RGB and thermal modalities while maintaining high image quality. In Fig.7 b, we visually present the Gaussian distributions of the original 3DGS method and our method with multimodal regularization.
@@ 6 Conclusion and Future Work
We are the first to implement thermal reconstruction based on 3DGS. We not only achieve simultaneous rendering of thermal and RGB images but also significantly improve the rendering quality of both color and thermal images. Additionally, we greatly reduce the model’s storage requirements. In the appendix, we discuss the limitations of this work and potential directions for future research.
Acknowledgments
This work was supported by the Key Project of National Nature Science Foundation of China (U22A2047), the ”Pioneer” and ”Leading Goose” R&\&&D Program of Zhejiang Province(2023C01044, 2024C01107, 2023C01030, 2023C03012), the National Nature Science Foundation of China (62301198, 62371173).